OpenStudy (anonymous):

Can someone help me please? The perimeter of a rectangle equals the sum of the lengths of its four sides. If the width of a rectangle is 9 inches and its perimeter is 42 inches, find the length of the rectangle.

OpenStudy (anonymous):

Okay, so the perimeter of the rectangle equals the sum of it's four sides. If there are two widths and two heights, and the perimeter is 42, it would be 2h+2w=42. If we know the rectangle is 9 inches wide, we can put that in for w in the equation. 2h+2(9)=42 2h+18=42. You can subtract the 18 from both sides and get 24 2h=24 Divide by 2 And you would get 12 as the height.

OpenStudy (anonymous):

Thank you so much